College of Computer & Information Science Spring 2005 Northeastern University Handout 1 CSU690: Algorithms and Data 5 January 2005

Questionnaire

Name: Email address: Major:

1. Have you taken CSU370 (Object-Oriented Design) or equivalent?

2. Have you taken CSU390 (Theory of Computation) or equivalent?

3. Are you familiar with the following data structures?

• Linked lists:

• Stacks and queues:

• Binary search trees:

• Balanced binary search trees (e.g., AVL trees, 2-3 trees, red-black trees, B-trees):

• Hash tables:

4. Which sorting algorithms have you learned? For each one, write “yes” if you also analyzed the running time and “no” if not.